OpenStudy (anonymous):

exhibit a bijection between N and the set of all even integers which are strictly less than 2013.

OpenStudy (helder_edwin):

is N the set of natural numbers here?

OpenStudy (helder_edwin):

r u there @aceecabalquinto ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry about that lol

OpenStudy (helder_edwin):

Let \[ \large \begin{align*} f:\mathbb{N}&\to\{x\in\mathbb{Z}:x\text{ is even and }x<2013\}\\ n&\mapsto -n+2013 \end{align*} \]

OpenStudy (helder_edwin):

i'll let u prove it is a bijection. it is easy.

OpenStudy (anonymous):

why is it -n+2012? is "n" any number?

OpenStudy (anonymous):

2013*

OpenStudy (helder_edwin):

f(1)=-1+2013=2012

OpenStudy (helder_edwin):

i made a mistake. wait please.

OpenStudy (anonymous):

ok

OpenStudy (helder_edwin):

it should be \[ \large f:n\mapsto -2n+2014 \]

OpenStudy (anonymous):

how come?

OpenStudy (helder_edwin):

the idea was this \[ \large 1\mapsto2012\qquad2\mapsto2010\qquad3\mapsto2008 \]

OpenStudy (anonymous):

oh ok. that makes sense. i'm just gonna try out the bijection right now

OpenStudy (helder_edwin):

go ahead.

OpenStudy (anonymous):

ok so i got the injective part alright. i'm kinda confused with the surjection part